OpenStudy (anonymous):

an equatiom is shown 5/12+?=1/2

OpenStudy (anonymous):

\[\frac{ 5 }{ 12 }+?=\frac{ 1 }{ 2 }\] \[?=\frac{ 1 }{ 2 }-\frac{ 5 }{ 12 }=\frac{ 6-5 }{ 12 }=\frac{ 1 }{ 12 }\]